# Systematic study of confinement induced effects on atomic electronic structure

### v1

@@ Abstract
We point out that although a litany of studies have been published on atoms in hard-wall confinement, they have either not been systematic, having only looked at select atoms and/or select electronic configurations, or they have not used robust numerical methods. To remedy the situation, we perform in this work a methodical study of atoms in hard-wall confinement with the HelFEM program, which employs the finite element method (FEM) that trivially implements the hard-wall potential, guarantees variational results, and allows easily finding the numerically exact solution.
We systematically examine low lying configurations of the H–Xe atoms and their monocations, and investigate how the states—especially the ground state—behave as a function of the position of the hard-wall boundary. Our fully numerical calculations are based on non-relativistic density functional theory (DFT), employing the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional and spherically averaged densities. We consider calculations with both spin-polarized as well as spin-restricted densities, and demonstrate that spin-polarization effects are significant in open shell configurations, even though some previous studies have only considered the spin-restricted model.
We demonstrate the importance of considering ground state changes for confined atoms by computing the ionization radii for the H–Xe atoms and observe significant differences to earlier studies. Confirming previous observations, we identify electron shifts on the outermost shells for a majority of the elements: valence $s$ electrons are highly unfavored under strong confinement, and the high-lying $3d$ and $4f$ orbitals become occupied in atoms of periods 2–3 and 4, respectively.
@@ 1 Introduction
The pioneering study of Michels et al.1 was concerned on the question of how atomic polarizabilities—which have an important role in chemical bonding—evolve as a function of pressure. Published in 1937, the study examined the hydrogen atom enclosed in an impenetrable sphere of a variable radius to simulate pressure effects. Such frugal models offer a great tool for understanding basic physical effects precisely due to their simplicity: Michels et al.1 found that hydrogen becomes less polarizable in increasing pressure.
@@
An interest in analogous studies of atoms in hard-wall confinement continues to this day. A multitude of studies have been published on the very first atoms of the periodic table; see ref. 2 for a review dedicated to the H and He atoms confined by a finite or infinite spherical barrier with 200 references. Yet, few studies have considered many-electron atoms, or reported systematic calculations on the periodic table.
@@
In hard-wall confinement, the wave function must vanish beyond the hard-wall boundary, and this can only be accomplished within a basis set with finite support. Although studies on confined atoms employing basis sets of an analytic form—such as Gaussian or Slater type orbital basis sets—have been published, they have finite tails in the forbidden region. Truncated Gaussian or Slater orbital basis sets do not have this issue, but they also suffer from basis set truncation errors. Our focus here is on fully numerical studies that are able to both accurately describe confinement effects as well as recover the numerically exact wave function.
@@
We will now briefly review the literature we found that fulfills our key criteria at least in part: studies mentioned herein have examined atoms with more than 4 electrons, and employed a numerically robust method.
@@
Boeyens3 performed fully numerical Hartree–Fock calculations to determine the ionization radii of compressed atoms. However, Boeyens3 did not employ a proper hard-wall potential. To approximate a hard-wall boundary at $r=r_{\mathrm{c}}$, Boeyens3 multiplied the wave function by the smooth step function $\Theta(r)=\exp\left[-\left(r/r_{c}\right)^{20}\right]$ at every self-consistent field (SCF) iteration, before the normalization of the updated wave function was carried out. Boeyens3 employed standard electronic configurations for the elements, and thus did not consider changes in the occupations as a function of the confinement.
@@
Sarkar et al.6 studied the effects of confinement on the chemical reactivity of the nitrogen atom. Sarkar et al.6 compared the method of Boeyens3 to the true hard-wall potential, and found the arising chemical reactivities to yield the same qualitative trends. Moreover, Sarkar et al.6 found that different methods to calculate the reactivity yield dissimilar results, and that different electronic occupations also lead to different chemical reactivities.
@@
Connerade et al.7 studied the filling of shells in the $3d$ and $4d$ elements as well as their cations in average-of-configurations Hartree–Fock calculations in a spherical cavity surrounded by a $10E_{h}$ potential wall. They found that orbital filling becomes more regular for successive rows with increasing pressure. In specific, the $s$$d$ competition disappears for these elements, as the $(n-1)d$ orbitals become energetically more favourable than the $ns$ orbitals under confinement. Connerade and coworkers have also reported studies on some individual atoms under confinement. For instance, Connerade and Dolmatov examined the Cr atom, while Connerade and Semaoune performed relativistic calculations on the La and Cs atoms in refs. 9 and 10, respectively.
@@
A number of studies of atoms in hard-wall confinement have been carried out at the DFT level by Garza et al.. Garza et al.11 reported a finite differences method (FDM) implementation for solving the atomic DFT equations in the presence of a hard-wall boundary, and studied the Ne and Na (and He) atoms in exchange-only calculations. In a follow-up study, Garza et al.12 examined the electrostatic potential of the unconfined and hard-wall confined Li, Na, K, and Rb atoms with DFT in order to study changes in the atoms’ shell structure, and found that shell structure is gradually lost in increasing confinement. Similarly, Sen et al.13 determined atomic ionization radii for the He–Ca atoms with hard-wall DFT calculations. As none of the studies by Garza et al.11, Garza et al.12, or Sen et al.13 discuss occupations, we can only assume that the ground state electron configuration of the unconfined atom was used in their calculations. In contrast, the study on the chemical reactivity indices of the compressed Li, Na, and K; Be, Mg, and Ca; N, P, and As; and Ne, Ar, and Kr atoms by Garza et al.14, which was likewise carried out with hard-wall DFT, explicitly mentions finding the lowest electronic configuration for each confinement radius.
@@
Continuing with the Garza group, Guerra et al.15 modeled pressure effects on the electronic properties of Ca, Sr, and Ba with hard-wall DFT, and found that their ground state changes from the $ns^{2}$ configuration to the $ns^{1}(n-1)d^{1}$ and $ns^{0}(n-1)d^{2}$ configurations in increasing pressure. Confinement effects on the spin potential of first-row transition metal cations were investigated at the hard-wall DFT level of theory by Lozano-Espinosa et al.16.
@@
Pašteka et al.17 studied the C and K (and He) atoms in hard confinement with the finite element method (FEM) at the complete active space self-consistent field18, 19 (CASSCF) level of theory, which represent the most sophisticated calculations in the studies we found. Also Pašteka et al.17 found changes in atomic ground state configurations under confinement: the ground state configuration of carbon changes from $1s^{2}2s^{2}2p^{2}$ in the free atom to $1s^{2}2p^{4}$ in strong confinement.
@@
As this literature overview demonstrates, although several fully numerical studies have been published on atoms in confinement, they have mostly focused on small groups of similar atoms. The study of Boeyens3 was the most complete, considering non-relativistic calculations on the first 105 atoms of the periodic table, but the study relied on an ad hoc implementation of confinement, and—like in many other studies mentioned above—changes in the ground state as a function of confinement were not considered. Studies that considered changes in the ground state are likewise limited: Connerade et al.7 only considered $3d$ and $4d$ elements, Garza et al.14 studied only atoms of groups 1, 2, 15, and 18, and Guerra et al.15 only studied heavy atoms of group 2.
@@
The only study that appears to have systematically investigated effects of confinement on the electronic structure of atoms so far is the study of Rahm et al.20, which employed the extreme pressure polarizable continuum model21 (XP-PCM) to study the first 93 atoms under pressure, employing Gaussian basis sets with scalar relativistic corrections and the Perdew–Burke–Ernzerhof (PBE) functional,22, 23 as well as its hybrid with 25% exact exchange, the PBE0 functional.24, 25 Rahm et al.20 identified electron shifts, $s\to p$, $s\to d$, $s\to f$, and $d\to f$, as essential chemical and physical consequences of compression.
As Gaussian basis sets cannot be expected to deliver reliable results in strong confinement,17 there is clearly a hole in the literature regarding systematic fully numerical studies covering (a significant fraction of) the periodic table.
@@
As is clear from the above discussion, changes in the electronic structure of atoms are an expected effect of confinement, and they thus need to be taken into account by considering changes in the electronic occupations as a function of confinement. In this work, we will address this major shortcoming of the existing literature with a systematic study on atoms in hard-wall confinement within a fully numerical approach, while also thoroughly considering the effect of the confinement on the ground state electron configuration.
@@
Our results confirm previous partial findings of the fully numerical studies of Connerade et al.7, Guerra et al.15, and Pašteka et al.17 as well as the thorough Gaussian basis study of Rahm et al.20: confinement induced electron shifts can be observed for a majority of the elements. It is important to note here that all of these studies have been carried out with different methodologies. This suggests that the findings of this work are indeed reliable despite the arguably simplistic level of theory employed herein.
@@
The layout of this work is the following. The form of the Hamiltonian including the hard-wall potential, as well as the level of theory is briefly discussed in section 2. The computational details including the employed numerical approach are presented in section 3. The results are presented in section 4. We first demonstrate that our method works by computing ionization energies and radii in section 4.1 and section 4.2 respectively. Then, we perform a systematic study of the behavior of various low lying configurations of the H–Xe atoms in hard-wall confinement. The full set of raw data as well as a detailed analysis of the results is included in the Supporting Information (SI); in section 4.3 we present the highlights from this analysis. The article concludes in a summary and discussion in section 5. Atomic units are employed throughout the work.
@@ 2 Theory
As already mentioned in section 1, we carry out calculations on many-electron atoms within the context of DFT.26, 27 We expand the one-particle states also known as atomic orbitals (AOs) in a numerical basis set as
@@
where $Y_{\mathrm{lm}}(\hat{\mathbf{r}})$ are spherical harmonics; we will discuss the form of the radial expansion $R^{\sigma}_{nl}(r)$ below in section 3.
Following standard practice in numerical DFT calculations on atoms, we assume that the electron density is spherically symmetric, $n_{\sigma}({\bf r})=n_{\sigma}(r)$ for each spin $\sigma$. This simplification leads to an efficient algorithm, since the problem of solving the Kohn–Sham equations,27 i.e., the optimization of the $s$, $p$, $d$, and $f$ orbitals becomes a set of coupled one-dimensional problems.28, 29
@@
$n_{\sigma}(r)=\sum_{nl}f^{\sigma}_{nl}[R^{\sigma}_{nl}(r)]^{2},$ (5)
where $f^{\sigma}_{nl}$ is the number of spin-$\sigma$ electrons on the $nl$ shell. We refer to refs. 28 and 29 for further discussion.
@@
The above choice of theory has both a physical and practical rationale. The fractional occupation formalism automatically guarantees spherical degeneracy,28, 29 since all magnetic sublevels are assumed to behave identically. In contrast, DFT calculations with integer occupations of the magnetic sublevels often break spatial symmetry and fail to reproduce the degeneracy of atomic multiplets.28 Although several specialized approaches to reproduce atomic multiplets within a DFT style description have been proposed,30, 31, 32, 33, 34 they appear to be quite involved and have not gained wide adoption. We also wish to direct the reader to the discussion of Baerends et al.35 on atomic reference energies for DFT calculations.

### v2

@@ Abstract
We point out that although a litany of studies have been published on atoms in hard-wall confinement, they have either not been systematic, having only looked at select atoms and/or select electron configurations, or they have not used robust numerical methods.
To remedy the situation, we perform in this work a methodical study of atoms in hard-wall confinement with the HelFEM program, which employs the finite element method (FEM) that trivially implements the hard-wall potential, guarantees variational results, and allows easily finding the numerically exact solution.
Our fully numerical calculations are based on non-relativistic density functional theory (DFT) and spherically averaged densities. We consider three levels of density functional approximations: the local density approximation employing the Perdew–Wang (PW92) functional, the generalized-gradient approximation (GGA) employing the Perdew–Burke–Ernzerhof (PBE) functional, and the meta-GGA approximation employing the r2SCAN functional. Importantly, the completely dissimilar density functional approximations are in excellent agreement, suggesting that the observed results are not artefacts of the employed level of theory.
We systematically examine low lying configurations of the H–Xe atoms and their monocations, and investigate how the configurations—especially the ground state configuration—behave as a function of the position of the hard-wall boundary. We perform calculations with both spin-polarized as well as spin-restricted densities, and demonstrate that spin-polarization effects are significant in open shell configurations, even though some previous studies have only considered the spin-restricted model.
We demonstrate the importance of considering ground state changes for confined atoms by computing the ionization radii for the H–Xe atoms and observe significant differences to earlier studies. Confirming previous observations, we identify electron shifts on the outermost shells for a majority of the elements: valence $s$ electrons are highly unfavored under strong confinement, and the high-lying $3d$ and $4f$ orbitals become occupied in atoms of periods 2–3 and 3–4, respectively.
We also comment on deficiencies of a commonly used density based estimate for the van der Waals (vdW) radius of atoms, and propose a better behaved variant in terms of the number of electrons outside the vdW radius that we expect will prove useful in future studies.
@@ 1 Introduction
The pioneering study of Michels et al.1 was concerned on the question of how atomic polarizabilities—which have an important role in chemical bonding—evolve as a function of pressure. Published in 1937, the study examined the hydrogen atom enclosed in an impenetrable sphere of a variable radius to simulate pressure effects. Such frugal models offer a great tool for understanding basic physical effects precisely due to their simplicity: Michels et al.1 found that hydrogen becomes less polarizable in increasing pressure, that is, when the radius of the sphere is decreased .
@@
An interest in analogous studies of atoms in hard-wall confinement (i.e. atoms in a spherical box) continues to this day. A multitude of studies have been published on the very first atoms of the periodic table; see ref. 2 for a review dedicated to the H and He atoms confined by a finite or infinite spherical barrier with 200 references. Yet, few studies have considered many-electron atoms, or reported systematic calculations on the periodic table.
@@
In hard-wall confinement, the wave function must vanish beyond the hard-wall boundary, and this can only be accomplished within a basis set with finite support. Many studies on confined atoms employing basis sets of an analytic form—such as Gaussian or Slater type orbital basis sets—have been published in the literature. However, these basis sets have finite tails in the forbidden region. Truncated Gaussian or Slater type orbital basis sets where this tail has been cut off do not have this issue, but they too suffer from basis set truncation errors (BSTEs), which are a priori unknown.
Our recent studies on atoms and molecules in strong magnetic fields found that changes in the electron configuration often resulted in significant increases in the BSTE of standard Gaussian-type orbital basis sets,3, 4 as these basis sets have not been optimized to describe electronic structure in this context. As we will discuss below, confinement is likewise expected to result in electron shifts to orbitals of higher angular momentum. Since this effect has again not been considered in the parametrization of standard Gaussian-type orbital basis sets, they cannot be expected to yield reliable estimates under confinement, since the orbitals of confined atoms can look extremely dissimilar to the orbitals of the unconfined atom. The qualitatively correct description of confinement induced effects then requires a careful study of basis set convergence, ideally supplemented by comparisons to fully numerical reference values. We point out that few studies have attempted to reach the complete basis set limit, and we will report numerically exact complete basis set limit energies later on in this work.
@@
Before laying out the scope of the present study, we will briefly review the literature we found that fulfills our key criteria at least in part: studies mentioned herein have examined atoms with more than 4 electrons, and employed a robust numerical method.
@@
Boeyens5 performed fully numerical Hartree–Fock calculations to determine the ionization radii of compressed atoms. However, Boeyens5 did not employ a proper hard-wall potential. To approximate a hard-wall boundary at $r=r_{\mathrm{c}}$, Boeyens5 multiplied the wave function by the smooth step function $\Theta(r)=\exp\left[-\left(r/r_{c}\right)^{20}\right]$ at every self-consistent field (SCF) iteration, before the normalization of the updated wave function was carried out. Boeyens5 employed standard electron configurations for the elements, and thus did not consider changes in the occupations as a function of the confinement.
@@
Sarkar et al.8 studied the effects of confinement on the chemical reactivity of the nitrogen atom. Sarkar et al.8 compared the method of Boeyens5 to the true hard-wall potential, and found the arising chemical reactivities to yield the same qualitative trends. Moreover, Sarkar et al.8 found that different methods to calculate the reactivity yield dissimilar results, and that different electronic occupations also lead to different chemical reactivities.
@@
Connerade et al.9 studied the filling of shells in the $3d$ and $4d$ elements as well as their cations in average-of-configurations Hartree–Fock calculations in a spherical cavity surrounded by a $10E_{h}$ potential wall. They found that orbital filling becomes more regular for successive rows with increasing pressure, that is, when the finite barrier is placed closer and closer to the nucleus. In specific, the $s$$d$ competition disappears for these elements, as the $(n-1)d$ orbitals become energetically more favourable than the $ns$ orbitals under confinement. Connerade and coworkers have also reported studies on some individual atoms under confinement. For instance, Connerade and Dolmatov10 examined the Cr atom, while Connerade and Semaoune performed relativistic calculations on the La and Cs atoms in refs. 11 and 12, respectively.
@@
A number of studies of atoms in hard-wall confinement have been carried out at the DFT level by Garza et al.. Garza et al.13 reported a finite differences method (FDM) implementation for solving the atomic DFT equations in the presence of a hard-wall boundary, and studied the Ne and Na (and He) atoms in exchange-only calculations. In a follow-up study, Garza et al.14 examined the electrostatic potential of the unconfined and hard-wall confined Li, Na, K, and Rb atoms with DFT in order to study changes in the atoms’ shell structure, and found that shell structure is gradually lost in increasing confinement. Similarly, Sen et al.15 determined atomic ionization radii for the He–Ca atoms with hard-wall confined DFT calculations. As none of the studies by Garza et al.13, Garza et al.14, or Sen et al.15 discuss occupations, we can only assume that the ground state electron configuration of the unconfined atom was used in their calculations. In contrast, the study on the chemical reactivity indices of the compressed Li, Na, and K; Be, Mg, and Ca; N, P, and As; and Ne, Ar, and Kr atoms by Garza et al.16, which was likewise carried out with hard-wall confined DFT, explicitly mentions finding the lowest electron configuration for each confinement radius.
@@
Continuing with the Garza group, Guerra et al.17 modeled pressure effects on the electronic properties of Ca, Sr, and Ba with hard-wall confined DFT, and found that their ground state changes from the $ns^{2}$ configuration to the $ns^{1}(n-1)d^{1}$ and $ns^{0}(n-1)d^{2}$ configurations in increasing pressure. Confinement effects on the spin potential of first-row transition metal cations were investigated at the hard-wall confined DFT level of theory by Lozano-Espinosa et al.18.
@@
Pašteka et al.19 studied the C and K (and He) atoms in hard-wall confinement with the finite element method (FEM) at the complete active space self-consistent field20, 21 (CASSCF) level of theory, which represents the most sophisticated level of theory in the studies we found. Also Pašteka et al.19 found changes in atomic ground state configurations under confinement: the ground state configuration of carbon changes from $1s^{2}2s^{2}2p^{2}$ in the free atom to $1s^{2}2p^{4}$ in strong confinement.
@@
As this literature overview demonstrates, although several fully numerical studies have been published on atoms in confinement, they have mostly focused on small groups of similar atoms. The study of Boeyens5 was the most complete, considering non-relativistic calculations on the first 105 atoms of the periodic table, but the study relied on an ad hoc implementation of confinement, and—like in many other studies mentioned above—changes in the ground state as a function of confinement were not considered. Studies that considered changes in the ground state are likewise limited: Connerade et al.9 only considered $3d$ and $4d$ elements, Garza et al.16 studied only atoms of groups 1, 2, 15, and 18, and Guerra et al.17 only studied heavy atoms of group 2.
@@
The only study that appears to have systematically investigated effects of confinement on the electronic structure of atoms so far is the study of Rahm et al.22, which employed the extreme pressure polarizable continuum model23 (XP-PCM) to study the first 93 atoms under pressure, employing the ANO-RCC Gaussian basis set24, *Roos2004_TCA_345, *Roos2004_JPCA_2851, *Roos2005_JPCA_9, *Roos2005_CPL_295, *Roos2008_JPCA_11431 with scalar relativistic corrections and the 25% hybrid30, 31 of the Perdew–Burke–Ernzerhof (PBE) functional,32, 33 which is commonly known as PBE0. Rahm et al.22 identified electron shifts, $s\to p$, $s\to d$, $s\to f$, and $d\to f$, as essential chemical and physical consequences of compression. In follow-up work, Rahm et al.34 studied non-bonded radii of compressed atoms with analogous calculations. Rahm et al.22 and Rahm et al.34 did not report carrying out a basis set convergence study of their findings. Yet, as already discussed above, Gaussian basis sets cannot be expected to deliver reliable results in strong confinement.19
As is clear from the above discussions, there is a hole in the literature regarding systematic fully numerical studies covering (a significant fraction of) the periodic table. Changes in the electronic structure of atoms are an expected effect of confinement, and they thus need to be taken into account by considering changes in the electronic occupations as a function of confinement. Fully numerical techniques are critially important as changes in the electronic occupations cannot be expected to be adequately reproduced by standard Gaussian-type orbital basis sets, for example.
@@
In this work, we will address this major shortcoming of the existing literature with a systematic study on atoms in hard-wall confinement within a fully numerical approach, while also thoroughly considering the effect of the confinement on the ground state electron configuration.
@@
Our results confirm previous partial findings of the fully numerical studies of Connerade et al.9, Guerra et al.17, and Pašteka et al.19 as well as the thorough Gaussian basis study of Rahm et al.22: confinement induced electron shifts can be observed for a majority of the elements. It is important to note here that all of these studies have been carried out with different methodologies. Furthermore, as we shall see, the dissimilar functionals we employ in this work are all in good agreement. This suggests that the findings of this work indeed correspond to physical effects of confinement despite the arguably simplistic level of theory employed herein.
@@
The layout of this work is the following. The theory behind the present work is briefly discussed in section 2. The form of the Hamiltonian including the hard-wall potential and the employed DFT approach is discussed in section 2.1. The employed theories for calculating ionization energies and ionization radii are discussed in sections 2.2 and 2.3, respectively. As our study includes data on a variety of electron configurations, we discuss ways to estimate the corresponding atomic radii in section 2.4, where we also introduce a new way to estimate atoms’ van der Waals (vdW) radii. Next, the computational details including the employed numerical approach are presented in section 3. The results are presented in section 4. We first demonstrate that our method works by computing ionization energies and ionization radii in sections 4.1 and 4.2, respectively. Then, we analyze the numerical and physical behavior of the employed vdW radius estimates in section 4.3 and find our vdW estimate to address many issues in a commonly used estimate. The highlights from our analysis of the behavior of the low lying configurations of the H–Xe atoms in hard-wall confinement are presented in section 4.4. The article concludes in a summary and discussion in section 5. Atomic units are employed throughout the work. The full set of raw data as well as a detailed analysis of the results is included in the Supporting Information (SI).
@@ 2.1 DFT approach
As already mentioned in section 1, we carry out calculations on many-electron atoms within the context of DFT.35, 36 We expand the one-particle states also known as atomic orbitals (AOs) in a numerical basis set as
@@
where $Y_{lm}(\hat{\mathbf{r}})$ are spherical harmonics; we will discuss the form of the radial expansion $R^{\sigma}_{nl}(r)$ below in section 3.
Following standard practice in numerical DFT calculations on atoms, we assume that the electron density is spherically symmetric, $n_{\sigma}({\bf r})=n_{\sigma}(r)$ for each spin $\sigma$. This simplification leads to an efficient algorithm, since the problem of solving the Kohn–Sham equations,36 i.e., the optimization of the $s$, $p$, $d$, and $f$ orbitals becomes a set of coupled one-dimensional problems.37, 38
@@
$n_{\sigma}(r)=\frac{1}{4\pi}\sum_{nl}f^{\sigma}_{nl}[R^{\sigma}_{nl}(r)]^{2},$ (5)
where $f^{\sigma}_{nl}$ is the number of spin-$\sigma$ electrons on the $nl$ shell. The occupations $f^{\sigma}_{nl}$ are determined from the number of electrons with angular momentum $l$ using Hund’s rules: the orbitals are filled starting from $n=l+1$ by occupying the $12l+$ 1 spin-up orbitals for the given $n$ before occupying the corresponding $12l+$ 1 spin-down orbitals. We refer to refs. 37 and 38 for further discussion.
@@
The above choice of theory has both a physical and a practical rationale. The fractional occupation formalism automatically guarantees spherical degeneracy,37, 38 since all magnetic sublevels are assumed to behave identically. In contrast, DFT calculations with integer occupations of the magnetic sublevels often break spatial symmetry and fail to reproduce the degeneracy of atomic multiplets.37 Although several specialized approaches to reproduce atomic multiplets within a DFT style description have been proposed,39, 40, 41, 42, 43 they appear to be quite involved and have not gained wide adoption. We also wish to direct the reader to the discussion of Baerends et al.44 on atomic reference energies for DFT calculations.
2.2 Ionization energies
Within the above level of theory, atoms’ ionization energies can be calculated in two ways. First, in the $\Delta$SCF method the ionization energy is determined for each value of $r_{c}$ as the difference between the total energy of the cation and of the neutral atom
$E_{+}(r_{c})=E_{\mathrm{cation}}(r_{c})-E_{\text{neutral atom}}(r_{c}).$ (6)
As is discussed below in section 3, the calculations on unconfined atoms employ the value $r_{c}=40a_{0}$ which is large enough to afford fully converged energies for the ground states of atoms and their cations.
Second, Boeyens5 used Koopmans’ theorem45 to determine the ionization radii. While there is no Koopmans’ theorem for DFT as in Hartree–Fock, Janak’s theorem46 provides a justification for an analysis based on orbital energies. The ionization energy is determined via Janak’s theorem as
$E_{+}(r_{c})=-\epsilon^{\mathrm{HOAO}}(r_{c}),$ (7)
where $\epsilon^{\mathrm{HOAO}}(r_{c})$ is the eigenvalue of the highest occupied atomic orbital (HOAO) of the neutral atom with the hard-wall boundary at $r=r_{c}$.
2.3 Ionization radii
Now that we have defined ways to calculate ionization energies, we can calculate ionization radii for atoms, i.e., the location of the hard-wall boundary where the atom favors casting away its outermost electron. The ionization radius $r_{+}$ for atom X is by solving the nonlinear equation
$E_{+}(r_{c})=0.$ (8)
If the ionization energy is computed with eq. 6, eq. 8 is straightforward to solve to high precision by bisection, once a crossing between the atomic and cationic curves has been identified. Now, the issue becomes that the energy of the atom in its neutral and cationic charge state depends on the employed electron configuration. This issue can be approached in two ways. The first is to follow the work of Boeyens5 and many others (see section 1) and fix the electron configuration to that of the ground state of the unconfined neutral atom and its monocation, respectively. The second is to realize that confinement can affect the ground state configurations of the atom and and its cation, and to compute the energies $E_{\text{neutral atom}}(r_{c})$ and $E_{\mathrm{cation}}(r_{c})$ for the lowest configurations of the neutral atom and cation, respectively, at each value of $r_{c}$.
Alternatively, we can use the estimate from Janak’s theorem of eq. 7 to solve eq. 8. Again, one can choose to use the electron configuration of the ground state of the unconfined neutral atom, or to employ a relaxed configuration at every value of $r_{c}$. However, exact solution to eq. 8 is sometimes problematic in the latter case. A case in point is the V atom: the value of $r_{+}$ corresponds to the location of a ground state crossing between a state with $\epsilon^{\mathrm{HOAO}}(r)<0$ and a state with $\epsilon^{\mathrm{HOAO}}(r)>0$, which means that eq. 8 is not truly satisfied no matter how tightly $r_{+}$ is converged.
The origin of the above problem is the present use of electron configurations with integer occupations. However, it has long been known that the use of non-integral occupation numbers for the $4s$ and $3d$ shells can result in a lower total energy for transition metal atoms in average-of-configurations Hartree–Fock calculations.47 This phenomenon is also topical for the present DFT calculations,48, 37 and has also been found in DFT calculations that do not employ any exchange-correlation functional.49 Calculations employing variationally optimized fractional occupation numbers would allow finding an exact solution to eq. 8, likely producing a slightly different radius. However, we do not consider such fractional occupations in this work, since integer occupation numbers are more commonly used and allow a clearer physical interpretation of the results.
2.4 Atomic radii
The location of the density maximum of the outermost orbital is a long-established estimator for the size of covalently bound atoms50
where $n_{i}(r)$ is the $i$th orbital density.
Since the electron density of an atom decays in the far valence region, a density threshold can be used to estimate the atom’s van der Waals (vdW) radius51
$n(r_{\rho})=\epsilon,$ (10)
where the electron density was defined above in eq. 5, and the solution of eq. 10 with a given threshold $\epsilon$ defines the vdW radius estimate $r_{\rho}$.
Due to several issues related to eq. 10 discovered and discussed in detail later in this work (for instance, eq. 10 does not have a solution in some cases, leaving the vdW radius undefined), we propose another metric for the vdW radius in this work. The vdW radius can be determined by containing all the electrons of the atom except $\epsilon$ within the volume enclosed by that radius,
$4\pi\int_{r_{\epsilon}}^{\infty}n(r)r^{2}\mathrm{d}r=\epsilon,$ (11)
where the electron density is again defined by eq. 5. This metric is well defined for any configuration of any element, and also exhibits superior numerical stability over eq. 10; see section 4.3 for discussion. Equation 11 is also straightforward to extend to non-spherically symmetric electron densities as
[…]
which will be relevant for applications with the CASSCF method, for instance.